# Time-Varying Gaussian Process Bandits with Unknown Prior

### v3

@@ 1 INTRODUCTION
Selecting prior by the maximising type-II likelihood can be a valid strategy in certain scenarios and statistical literature provides some asymptotic guarantees (Kaufman & Shaby, 2013; Li, 2022) as the number of observations increases, but they do not apply to all priors, particularly non-stationary ones. To universally guarantee that type-II MLE will recover the true prior (the prior black-box function was sampled from), one needs to cover the function domain with a dense, uniform grid (Bachoc, 2013). However, this defeats the purpose of sample efficient optimisation, where instead of querying the function everywhere, we wish to use our limited query budget in the region close to the optimal solution. We would thus want to have an algorithm, which conducts optimisation while simultaneously trying out different priors and selecting the most promising one.
If the black-box function and search space do not change over time, one may employ the Regret Balancing (Pacchiano et al., 2020; Ziomek et al., 2024) technique as a solution. Such a technique tries out multiple priors, analyses the function values obtained while using each of them and eliminates the underperforming ones. The regret of such a procedure is guaranteed to be ‘close’ to the regret of an algorithm using the best prior among the candidates. However, if the problem exhibits any form of time-varying behaviour, the difference in function values observed while using different priors does not necessarily be due to the choice of prior, but rather to changes in the underlying problem structure.
@@ 2 PROBLEM STATEMENT
We consider a problem setting where we wish to maximise an unknown, black-box and spatio-temporal function $f:\mathcal{X}\times[T]\to\mathbb{R}$ over some compact set $\mathcal{X}\subset\mathbb{R}^{d}$. We assume we are allowed to query a single point $\bm{x}_{t}$ at each timestep $t\in[T]$ and obtain its corrupted function value $y_{t}=f(\bm{x}_{t},t)+\epsilon_{t}$ by some i.i.d. Gaussian noise $\epsilon_{t}\sim\mathcal{N}(0,R^{2})$ for all $t\in[T]$. The function is allowed to change with time step $t$, hence the explicit dependence on $t$ in its signature. We also assume the available domain at each time step $\mathcal{X}_{t}\subseteq\mathcal{X}$ can change in an arbitrary manner, possibly chosen by an adversary. As such, formally, at each time step, the optimal point we can select is $\bm{x}^{*}_{t}=\arg\max_{\bm{x}\in\mathcal{X}_{t}}f(\bm{x},t)$. As we want to maximise the function, the instantaneous regret is defined as $r_{t}=f(\bm{x}^{*}_{t},t)-f(\bm{x}_{t},t)$. Similarly, we define cumulative regret as $R_{T}=\sum_{t=1}^{T}r_{t}$. We assume that the black-box function and its evolution through time are governed by some prior, i.e., $f\sim p^{\star}(f)$ and we refer to $p^{\star}(\cdot)$ as the true prior. We also assume we are given access to a set of priors $\mathcal{U}$ provided to us (e.g. by domain experts) that contains the true prior $p^{*}(f)$. However, we also assume that we do not know, which prior in $\mathcal{U}$ is the true one.
@@
Each GP prior $p(f)$ is uniquely defined by its spatio-temporal mean function $\mu^{p}(\cdot):\mathcal{X}\times[T]\to\mathbb{R}$ and kernel function $k^{p}(\cdot,\cdot):(\mathcal{X}\times[T])\times(\mathcal{X}\times[T])\to\mathbb{R}$. We require $\forall_{t\in[T]}\sup_{\bm{x}\in\mathcal{X}}|\mu^{p}(\bm{x},t)|<\infty$ for each $p\in\mathcal{U}$. Without the loss of generality, we also assume that $k^{p}(\cdot,\cdot)\leq 1$ for each $p\in\mathcal{U}$. An example of such spatio-temporal kernel could be the time-varying RBF kernel defined as:
@@
$k((\bm{x},t),(\bm{x}^{\prime},t^{\prime}))=\exp\left(-\frac{1}{2l}\lVert\bm{x}-\bm{x}^{\prime}\rVert_{2}^{2}\right)(1-\epsilon)^{|t-t^{\prime}|/2},$
@@
where $l$ and $\epsilon$ are hyperparameters. Note that one could think about spatio-temporal kernels as standard stationary kernels with the input space extended to include one additional variable $t$ (Van Vaerenbergh et al., 2012a, b). The only difference is that when conducting optimisation, we have no control over the choice of the $t$ variable, as opposed to the $x$ variable.
@@
As the difficulty of the problem depends heavily on the kernel $k^{p^{*}}(\cdot,\cdot)$ of the prior $p^{*}(\cdot)$ the function was sampled from, typically in preceding literature, the regret bound is expressed in terms of maximum information gain (MIG, Srinivas et al. (2010)) of kernel $k^{p}$, defined as $\gamma_{T}^{p}:=\max_{x_{1},\dots,x_{T}}\frac{1}{2}\log\lvert I+R^{-2}\bm{K}^{p}_{T}\rvert$, where $\mathbf{K}^{p}_{T}\in\mathbb{R}^{T\times T}$ with entries $(\mathbf{K}^{p}_{T})_{i,j}=k^{p}((\bm{x}_{i},i),(\bm{x}_{j},j))$. Srinivas et al. (2010) derived bounds on MIG for standard kernels and then Bogunovic et al. (2016) extended these bounds to spatio-temporal kernels.
@@
Gaussian Process Model In GP-based BO, at each timestep $t$, we contruct a surrogate model of the black-box function, based on the queries so far. Let $\mathcal{D}_{t-1}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{t-1}$ be the collection of all queried points and their corresponding noisy values up to timestep $t-1$. If we use a GP model with a prior $p(\cdot)$ then given the data so far $\mathcal{D}_{t-1}$, the model returns us the following mean $\mu^{p}_{t-1}(\bm{x})$ and variance $(\sigma^{p}_{t-1})^{2}(\bm{x})$ functions:
$\mu^{p}_{t-1}(\bm{x},t^{\prime})=$ $(\bm{k}^{p}_{t-1}(\bm{x},t^{\prime})^{T}(\mathbf{K}^{p}_{t-1}+R^{2}\mathbf{I})^{-1}$
$(\bm{y}_{t-1}-\mu^{p}(\bm{x},t^{\prime})))+\mu^{p}(\bm{x},t^{\prime})$
$(\sigma^{p}_{t-1})^{2}(\bm{x},t^{\prime})=$ $k^{p}((\bm{x},t^{\prime}),(\bm{x},t^{\prime}))-$
$\bm{k}^{p}_{t-1}(\bm{x},t^{\prime})^{T}(\mathbf{K}^{p}_{t-1}+R^{2}\mathbf{I})^{-1}\bm{k}^{p}_{t-1}(\bm{x},t^{\prime})$
@@
where $\bm{y}_{t-1}\in\mathbb{R}^{t-1}$ with elements $(\bm{y}_{t-1})_{i}=y_{i}$, $\bm{k}^{p}_{t-1}(\bm{x},t^{\prime})\in\mathbb{R}^{t-1}$ with elements $\bm{k}^{p}_{t-1}(\bm{x},t^{\prime})_{i}=k^{p}((\bm{x},t^{\prime}),(\bm{x}_{i},i))$. Note the distinction between $t$ and $t^{\prime}$. The former refers to the time step in optimisation, whereas the latter to the point in time for which we wish to make a prediction. While in the algorithm we will always use $t=t^{\prime}$, as we wish to make a prediction about current timestep, this distinction will be useful for proving one of suplementary Lemmas. While using mean and variance functions provided by the GP, we get the guarantee on predictive performance, as stated in the next Section.
@@ 3 PRELIMINARIES
We now recall well-known results from the literature. Those results would be later needed to prove the regret bound of our proposed algorithm. The original versions come from a stationary setting, however, due to the previously mentioned equivalence of spatiotemporal kernels and kernels with expanded input space, the results apply to our setting as well when $\mathcal{X}$ is replaced with $\mathcal{X}\times[T]$.
Theorem 3.1 (Lemma 5.1 of Srinivas et al. (2010)).
Let $f\sim p(f)$, and set $\beta^{p}_{t}=\sqrt{2\log\left(\frac{|\mathcal{X}|T\pi^{2}t^{2}}{6\delta_{A}}\right)}$. Then, with probability at least $1-\delta_{A}$, for all $\bm{x}\in\mathcal{X}$ and all $t,t^{\prime}\in[T]$:
$\left|f(\bm{x},t^{\prime})-\mu_{t-1}^{p}(\bm{x},t^{\prime})\right|\leq\beta^{p}_{t}\sigma_{t-1}^{p}(\bm{x},t^{\prime}).$
Note that the confidence parameter $\beta_{n}$ in Theorem 3.1, depends on $|\mathcal{X}|$ in the Bayesian setting and only makes sense if $|\mathcal{X}|<\infty$. If this is not the case, we need to adopt a standard assumption Srinivas et al. (2010).
@@
$P\left(\max_{t\in[T]}\sup_{x\in\mathcal{X}}\left\lvert\frac{\partial f}{\partial x_{j}}\right\rvert>L\right)\leq a^{p}\exp\left(-\frac{L^{2}}{b^{p}}\right),$
@@
Under that assumption, we can derive high-probability confidence intervals even if $|\mathcal{X}|=\infty$. Following Srinivas et al. (2010), purely for the sake of analysis, let us define a discretization $D_{t}$ of size $(\tau_{t})^{d}$ so that for all $x\in D_{t}$, $\lVert x-[x]_{t}\rVert_{1}\leq\frac{r}{d\tau_{t}}$, where $[x]_{t}$ denotes the closest point in $D_{t}$ to $x$. We now get the following Theorem.
Theorem 3.3 (Lemma 5.5 and 5.7 of Srinivas et al. (2010)).
Let Assumption 3.2 hold, set $\beta_{t}^{p}=\sqrt{2\log\left(\frac{\pi^{2}t^{2}}{3\delta_{A}}\right)+2d\log\left(Tdt^{2}rb^{p}\sqrt{\log\left(\frac{2da^{p}}{\delta_{A}}\right)}\right)}$ and set the size of discretisation to $\tau_{t}=dt^{2}rb^{p}\sqrt{\log\left(\frac{2da^{p}}{\delta_{A}}\right)}$. Then with probability at least $1-\delta_{A}$ we have $\forall\bm{x}\in\mathcal{X}$ and $\forall t,t^{\prime}\in[T]$:
$|f(\bm{x},t^{\prime})-\mu_{t-1}^{p}([\bm{x}]_{t},t^{\prime})|\leq\beta_{t}^{p}\sigma^{p}_{t-1}([\bm{x}]_{t},t^{\prime})+\frac{1}{t^{2}}.$
And for any $\bm{x}_{t}$ deterministic conditioned on $\mathcal{D}_{t-1}$ :
$|f(\bm{x}_{t},t)-\mu_{t-1}^{p}(\bm{x}_{t},t)|\leq\beta_{t}^{p}\sigma^{p}_{t-1}(\bm{x}_{t},t).$
@@ 4 GP-UCB WITH PRIOR ELIMINATION
We now state our proposed algorithm in Algorithm 1. When selecting the next point to query (line 7), our algorithm is doubly optimistic, with respect to both the function value and the model. In other words, for each point, we use the largest possible upper confidence bound among the ones suggested by models defined by possible priors in $\mathcal{U}_{t-1}$. We denote the selected point by $\bm{x}_{t}$ and the prior under which it achieved the highest UCB value is denoted by $p_{t}(\cdot)$. The set $S^{p}_{t}$ stores all iterations $i$ up to the current one $t$, during which $p_{t}(\cdot)=p(\cdot)$, for each $p\in\mathcal{U}$ (line 8). After observing the function value, we calculate the error $\eta_{t}$ between the observed value and the mean at point $\bm{x}_{t}$, according to the model defined by $p_{t}(\cdot)$ (line 11). Then, each model for which the sum of errors is too big is removed from the set of candidates (lines 12-14) and we continue with the loop.
The intuition behind this mechanism is depicted in Figure 1. Let us assume we have two models GP1 and GP2, represented by red and blue solid lines, together with their corresponding confidence intervals in dashed lines. After we query a new point, it can either lie within the confidence intervals of both models (like the star) or only one model (like the diamond). If we were to obtain a point like the diamond, we clearly see that the red model’s prediction was very inaccurate and we lean towards rejecting the prior that produced it. On the other hand, if we observe a point like the star when querying the function then it does not matter, which model is the ”true” one, as such an observation is probable under both models and the errors for either of the models would be small. Thanks to this mechanism, we can conduct BO, whilst simultaneously eliminating the (most-likely) wrong priors. For such a strategy, we can derive a regret bound, as stated by the following Theorem.
@@
4:        Fit a GP $(\mu^{p}_{t-1}(\bm{x},t),\sigma^{p}_{t-1}(\bm{x},t))$ to $\mathcal{D}_{t-1}$
5:        Set $\textrm{UCB}_{t-1}^{p}(\bm{x})=\mu^{p}_{t-1}(\bm{x},t)+\beta_{t}^{p}\sigma^{p}_{t-1}(\bm{x},t)$
@@
7:     Find $\bm{x}_{t},p_{t}=\underset{\bm{x},p\in\mathcal{X}\times\mathcal{U}_{t-1}}{\arg\max}\textrm{UCB}_{t-1}^{p}(\bm{x})$
@@
11:     Set $\eta_{t}=y_{t}-\mu^{p_{t}}_{t-1}(\bm{x}_{t})$
12:     if $\left|\sum_{i\in S^{p_{t}}_{t}}\eta_{i}\right|>\sqrt{\xi_{t}|S^{p_{t}}_{t}|}+\sum_{i\in S^{p_{t}}_{t}}\beta^{p_{t}}_{i}\sigma_{i}^{p_{t}}(\bm{x}_{i})$ then
@@
Let $f\sim p^{*}(f)$, $\epsilon_{t}\sim\mathcal{N}(0,R^{2})$ for each $t$ and $p^{*}\in\mathcal{U}$ and denote $\gamma_{T}=\max_{p\in\mathcal{U}}\gamma_{T}^{p}$. Set $\xi_{t}=2R^{2}\log\frac{|\mathcal{U}|\pi^{2}t^{2}}{3\delta}$ and :
@@
    if $|\mathcal{X}|<\infty$, set $\beta_{t}^{p}=\sqrt{2\log\frac{T|\mathcal{X}|\pi^{2}t^{2}}{3\delta}}$
@@
$\beta_{t}^{p}=\sqrt{2\log\left(\frac{\pi^{2}t^{2}}{3\delta_{A}}\right)+2d\log\left(Tdt^{2}rb^{p}\sqrt{\log\left(\frac{2da^{p}}{\delta_{A}}\right)}\right)}$
@@
$R_{T}\leq\mathcal{O}\left(|\mathcal{U}|B+\sqrt{|\mathcal{U}|\xi_{T}T}+\beta_{T}\sqrt{|\mathcal{U}|\gamma_{T}T}\right)$
for some constant $B>0$.

### v4

@@ 1 INTRODUCTION
Selecting prior by the maximising type-II likelihood (Good, 1983; Berger, 1985) can be a valid strategy in certain scenarios and statistical literature provides some asymptotic guarantees (Kaufman & Shaby, 2013; Li, 2022) as the number of observations increases, but they do not apply to all priors, particularly non-stationary ones. To universally guarantee that type-II MLE will recover the true prior (the prior black-box function was sampled from), one needs to cover the function domain with a dense, uniform grid (Bachoc, 2013). However, this defeats the purpose of sample efficient optimisation, where instead of querying the function everywhere, we wish to use our limited query budget in the region close to the optimal solution. We would thus want to have an algorithm, which conducts optimisation while simultaneously trying out different priors and selecting the most promising one.
If the black-box function and search space do not change over time, one may employ the Regret Balancing (Pacchiano et al., 2020; Ziomek et al., 2024) technique as a solution. Such a technique tries out multiple priors, analyses the function values obtained while using each of them and eliminates the underperforming ones. The regret of such a procedure is guaranteed to be ‘close’ to the regret of an algorithm using the best prior among the candidates. However, if the problem exhibits any form of time-varying behaviour, the difference in function values observed while using different priors does not necessarily be due to the choice of prior, but rather to changes in the underlying problem.
@@
Related Work on BO with Misspecification The work of Bogunovic & Krause (2021) studied BO under model misspecification from the frequentist perspective, however, provided no method for finding the well-specified model. Hvarfner et al. (2023) proposed an algorithm that combines BO with active learning so that prior hyperparameters can be learnt simultanously while conducting optimisation. However, their method lacks theoretical guarantees and introduces substantial computational overhead. Rodemann & Augustin (2024) proposed an algorithm to conduct BO when instead of one prior mean function, we are provided a set of possible mean functions, however, provided no regret guarantees for this algorithm. Xu et al. (2024) proposed an algorithm for conducting BO with preferential feedback, where the black-box function value itself cannot be directly evaluated. In a way, one can think about this problem settings as a misspecification of the observed function value. To address this authors propose to use likelihood based confidence sets and remain optimistic with respect to the best feasible function in the confidence set. However, this work assumes frequentist setting and that kernel hyperparameters are known.
@@ 2 PROBLEM STATEMENT
We consider a problem setting where we wish to maximise an unknown, black-box and spatio-temporal function $f:\mathcal{X}\times[T]\to\mathbb{R}$ over some compact set $\mathcal{X}\subset\mathbb{R}^{d}$. We assume we are allowed to query a single point $\mathbf{x}_{t}$ at each timestep $t\in[T]$ and obtain its corrupted function value $y_{t}=f(\mathbf{x}_{t},t)+\epsilon_{t}$ by some i.i.d. Gaussian noise $\epsilon_{t}\sim\mathcal{N}(0,R^{2})$ for all $t\in[T]$. The function is allowed to change with time step $t$, hence the explicit dependence on $t$ in its signature. We also assume the available domain at each time step $\mathcal{X}_{t}\subseteq\mathcal{X}$ can change in an arbitrary manner, possibly chosen by an adversary. As such, formally, at each time step, the optimal point we can select is $\mathbf{x}^{*}_{t}=\arg\max_{\mathbf{x}\in\mathcal{X}_{t}}f(\mathbf{x},t)$. As we want to maximise the function, the instantaneous regret is defined as $r_{t}=f(\mathbf{x}^{*}_{t},t)-f(\mathbf{x}_{t},t)$. Similarly, we define cumulative regret as $R_{T}=\sum_{t=1}^{T}r_{t}$. We assume that the black-box function and its evolution through time are governed by some prior, i.e., $f\sim p^{\star}(f)$ and we refer to $p^{\star}(\cdot)$ as the true prior. We also assume we are given access to a set of priors $\mathcal{U}$ provided to us (e.g. by domain experts) that contains the true prior $p^{*}(f)$. However, we also assume that we do not know, which prior in $\mathcal{U}$ is the true one.
@@
Each GP prior $p(f)$ is uniquely defined by its spatio-temporal mean function $\mu^{p}(\cdot):\mathcal{X}\times[T]\to\mathbb{R}$ and kernel function $k^{p}(\cdot,\cdot):(\mathcal{X}\times[T])\times(\mathcal{X}\times[T])\to\mathbb{R}$. We require $\forall_{t\in[T]}\sup_{\mathbf{x}\in\mathcal{X}}|\mu^{p}(\mathbf{x},t)|<\infty$ for each $p\in\mathcal{U}$. Without the loss of generality, we also assume that $k^{p}(\cdot,\cdot)\leq 1$ for each $p\in\mathcal{U}$. An example of such spatio-temporal kernel could be the time-varying RBF kernel:
@@
$k((\mathbf{x},t),(\mathbf{x}^{\prime},t^{\prime}))=\exp\left(-\frac{1}{2l}\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert_{2}^{2}\right)(1-\varepsilon)^{|t-t^{\prime}|/2},$
@@
where $l$ and $\varepsilon$ are hyperparameters. Note that one could think about spatio-temporal kernels as standard stationary kernels with the input space extended to include one additional variable $t$ (Van Vaerenbergh et al., 2012a, b). The only difference is that when conducting optimisation, we have no control over the choice of the $t$ variable, as opposed to the $\mathbf{x}$ variable.
@@
As the difficulty of the problem depends heavily on the kernel $k^{p^{*}}(\cdot,\cdot)$ of the prior $p^{*}(\cdot)$ the function was sampled from, typically in preceding literature, the regret bound is expressed in terms of maximum information gain (MIG, Srinivas et al. (2010)) of kernel $k^{p}$, defined as $\gamma_{T}^{p}:=\max_{x_{1},\dots,x_{T}}\frac{1}{2}\log\lvert I+R^{-2}\mathbf{K}^{p}_{T}\rvert$, where $\mathbf{K}^{p}_{T}\in\mathbb{R}^{T\times T}$ with entries $(\mathbf{K}^{p}_{T})_{i,j}=k^{p}((\mathbf{x}_{i},i),(\mathbf{x}_{j},j))$. Srinivas et al. (2010) derived bounds on MIG for standard kernels and then Bogunovic et al. (2016) extended these bounds to spatio-temporal kernels.
@@
Gaussian Process Model In GP-based BO, at each timestep $t$, we contruct a surrogate model of the black-box function, based on the queries so far. Let $\mathcal{D}_{t-1}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{t-1}$ be the collection of all queried points and their corresponding noisy values up to timestep $t-1$. If we use a GP model with a prior $p(\cdot)$ then given the data so far $\mathcal{D}_{t-1}$, the model returns us the following mean $\mu^{p}_{t-1}(\mathbf{x})$ and variance $(\sigma^{p}_{t-1})^{2}(\mathbf{x})$ functions:
$\mu^{p}_{t-1}(\mathbf{x},t^{\prime})=$ $(\mathbf{k}^{p}_{t-1}(\mathbf{x},t^{\prime})^{T}(\mathbf{K}^{p}_{t-1}+R^{2}\mathbf{I})^{-1}$
$(\mathbf{y}_{t-1}-\mu^{p}(\mathbf{x},t^{\prime})))+\mu^{p}(\mathbf{x},t^{\prime})$
$(\sigma^{p}_{t-1})^{2}(\mathbf{x},t^{\prime})=$ $k^{p}((\mathbf{x},t^{\prime}),(\mathbf{x},t^{\prime}))-$
$\mathbf{k}^{p}_{t-1}(\mathbf{x},t^{\prime})^{T}(\mathbf{K}^{p}_{t-1}+R^{2}\mathbf{I})^{-1}\mathbf{k}^{p}_{t-1}(\mathbf{x},t^{\prime})$
@@
where $\mathbf{y}_{t-1}\in\mathbb{R}^{t-1}$ with elements $(\mathbf{y}_{t-1})_{i}=y_{i}$, $\mathbf{k}^{p}_{t-1}(\mathbf{x},t^{\prime})\in\mathbb{R}^{t-1}$ with elements $\mathbf{k}^{p}_{t-1}(\mathbf{x},t^{\prime})_{i}=k^{p}((\mathbf{x},t^{\prime}),(\mathbf{x}_{i},i))$. Note the distinction between $t$ and $t^{\prime}$. The former refers to the time step in optimisation, whereas the latter to the point in time for which we wish to make a prediction. While in the algorithm we will always use $t=t^{\prime}$, as we wish to make a prediction about current timestep, this distinction will be useful for proving one of suplementary Lemmas. While using mean and variance functions provided by the GP, we get the guarantee on predictive performance, as stated in the next Section.
@@ 3 PRELIMINARIES
We now recall well-known results from the literature. Those results would be later needed to prove the regret bound of our proposed algorithm. This results come from the work of Bogunovic et al. (2016), which extend the stationary setting result from Srinivas et al. (2010).
Theorem 3.1 (Appendix C1 of Bogunovic et al. (2016)).
Let $f\sim p(f)$, and set $\beta^{p}_{t}=\sqrt{2\log\left(\frac{|\mathcal{X}|\pi^{2}t^{2}}{2\delta_{A}}\right)}$. Then, with probability at least $1-\delta_{A}$, for all $\mathbf{x}\in\mathcal{X}$ and all $t\in[T]$:
$\left|f(\mathbf{x},t)-\mu_{t-1}^{p}(\mathbf{x},t)\right|\leq\beta^{p}_{t}\sigma_{t-1}^{p}(\mathbf{x},t).$
Note that the confidence parameter $\beta_{t}^{p}$ in Theorem 3.1, depends on $|\mathcal{X}|$ in the Bayesian setting and only makes sense if $|\mathcal{X}|<\infty$. If this is not the case, we need to adopt a standard assumption (Srinivas et al., 2010; Bogunovic et al., 2016).
@@
$P\left(\sup_{x\in\mathcal{X}}\left\lvert\frac{\partial f}{\partial x_{j}}\right\rvert>L\right)\leq a^{p}\exp\left(-\frac{L^{2}}{b^{p}}\right),$
@@
Under that assumption, we can derive high-probability confidence intervals even if $|\mathcal{X}|=\infty$. Following (Srinivas et al., 2010; Bogunovic et al., 2016), purely for the sake of analysis, let us define a discretization $D_{t}$ of size $(\tau_{t})^{d}$ so that for all $x\in D_{t}$, $\lVert x-[x]_{t}\rVert_{1}\leq\frac{r}{d\tau_{t}}$, where $[x]_{t}$ denotes the closest point in $D_{t}$ to $x$. We now get the following Theorem.
Theorem 3.3 (Appendix C1 of Bogunovic et al. (2016)).
Let Assumption 3.2 hold, set $\beta_{t}^{p}=\sqrt{2\log\left(\frac{\pi^{2}t^{2}}{2\delta_{A}}\right)+2d\log\left(dt^{2}rb^{p}\sqrt{\log\left(\frac{da^{p}}{2\delta_{A}}\right)}\right)}$ and set the size of discretisation to $\tau_{t}=dt^{2}rb^{p}\sqrt{\log\left(\frac{2da^{p}}{\delta_{A}}\right)}$. Then with probability at least $1-\delta_{A}$ we have $\forall\mathbf{x}\in\mathcal{X}$ and $\forall t\in[T]$:
$|f(\mathbf{x},t)-\mu_{t-1}^{p}([\mathbf{x}]_{t},t)|\leq\beta_{t}^{p}\sigma^{p}_{t-1}([\mathbf{x}]_{t},t)+\frac{1}{t^{2}}.$
And for any $\mathbf{x}_{t}$ deterministic conditioned on $\mathcal{D}_{t-1}$ :
$|f(\mathbf{x}_{t},t)-\mu_{t-1}^{p}(\mathbf{x}_{t},t)|\leq\beta_{t}^{p}\sigma^{p}_{t-1}(\mathbf{x}_{t},t).$
@@ 4 GP-UCB WITH PRIOR ELIMINATION
We now state our proposed algorithm in Algorithm 1. When selecting the next point to query (line 7), our algorithm is doubly optimistic, with respect to both the function value and the model. In other words, for each point, we use the largest possible upper confidence bound among the ones suggested by models defined by possible priors in $\mathcal{U}_{t-1}$. We denote the selected point by $\mathbf{x}_{t}$ and the prior under which it achieved the highest UCB value is denoted by $p_{t}(\cdot)$. The set $S^{p}_{t}$ stores all iterations $i$ up to the current one $t$, during which $p_{t}(\cdot)=p(\cdot)$, for each $p\in\mathcal{U}$ (line 8). After observing the function value, we calculate the error $\eta_{t}$ between the observed value and the mean at point $\mathbf{x}_{t}$, according to the model defined by $p_{t}(\cdot)$ (line 11). Then, each model for which the sum of errors is too big is removed from the set of candidates (lines 12-14) and we continue with the loop. As we will show later, the true prior $p^{\star}$ will never be rejected with high probability. As such, if all priors have been rejected, this likely indicates that none of the supplied priors was the true prior. In such case, the algorithm could terminate and flag this issue to the user.
The intuition behind the rejection mechanism is depicted in Figure 1. Let us assume we have two models GP1 and GP2, represented by red and blue solid lines, together with their corresponding confidence intervals in dashed lines. After we query a new point, it can either lie within the confidence intervals of both models (like the star) or only one model (like the diamond). If we were to obtain a point like the diamond, we clearly see that the red model’s prediction was very inaccurate and we lean towards rejecting the prior that produced it. On the other hand, if we observe a point like the star when querying the function then it does not matter, which model is the "true" one, as such an observation is probable under both models and the errors for either of the models would be small. Thanks to this mechanism, we can conduct BO, whilst simultaneously eliminating the (most-likely) wrong priors. For such a strategy, we can derive a regret bound, as stated by the following Theorem.
@@
4:        Fit a GP $(\mu^{p}_{t-1}(\mathbf{x},t),\sigma^{p}_{t-1}(\mathbf{x},t))$ to $\mathcal{D}_{t-1}$
5:        Set $\textrm{UCB}_{t-1}^{p}(\mathbf{x})=\mu^{p}_{t-1}(\mathbf{x},t)+\beta_{t}^{p}\sigma^{p}_{t-1}(\mathbf{x},t)$
@@
7:     Find $\mathbf{x}_{t},p_{t}=\underset{\mathbf{x},p\in\mathcal{X}\times\mathcal{U}_{t-1}}{\arg\max}\textrm{UCB}_{t-1}^{p}(\mathbf{x})$
@@
11:     Set $\eta_{t}=y_{t}-\mu^{p_{t}}_{t-1}(\mathbf{x}_{t})$
12:     if $\left|\sum_{i\in S^{p_{t}}_{t}}\eta_{i}\right|>\sqrt{\xi_{t}|S^{p_{t}}_{t}|}+\sum_{i\in S^{p_{t}}_{t}}\beta^{p_{t}}_{i}\sigma_{i}^{p_{t}}(\mathbf{x}_{i})$ then
@@
Let $f\sim p^{*}(f)$, $\epsilon_{t}\sim\mathcal{N}(0,R^{2})$ for each $t$ and $p^{*}\in\mathcal{U}$ and denote $\gamma_{T}=\max_{p\in\mathcal{U}}\gamma_{T}^{p}$. Set $\xi_{t}=2R^{2}\log\frac{|\mathcal{U}|\pi^{2}t^{2}}{\delta}$ and :
@@
    if $|\mathcal{X}|<\infty$, set $\beta_{t}^{p}=\sqrt{2\log\frac{2|\mathcal{X}|\pi^{2}t^{2}}{\delta}}$
@@
$\beta_{t}^{p}=\sqrt{2\log\left(\frac{3\pi^{2}t^{2}}{2\delta}\right)+2d\log\left(dt^{2}rb^{p}\sqrt{\log\left(\frac{3da^{p}}{2\delta}\right)}\right)}$
@@
$R_{T}\leq\mathcal{O}\left(|\mathcal{U}|B_{t^{\star}}+\sqrt{|\mathcal{U}|\xi_{T}T}+\beta_{T}\sqrt{|\mathcal{U}|T\gamma}_{T}\right)$
where $\beta_{T}=\max_{p\in\mathcal{U}}\beta_{T}^{p}$, $\gamma_{T}=\max_{p\in\mathcal{U}}\gamma_{T}^{p}$ and $t^{\star}$ is the last iteration where any prior was rejected and $B_{t^{\star}}=\beta_{t}^{p^{\star}}+\max_{t^{\prime}\in[t^{\star}]}\sup_{\mathbf{x}\in\mathcal{X}}|\mu^{p^{\star}}(\mathbf{x},t^{\prime})|+1$.